OpenStudy (anonymous):

Write the equation of the line that goes through the points (using the given form) (0,5) and (0,6), standard NEED QUICK HELP

OpenStudy (john_es):

Are you sure these are the points?

OpenStudy (anonymous):

Oh I'm sorry (6,0)

OpenStudy (john_es):

So, (0,5) and (6,0). I suppose the standard form of the line is this, \[y=m\cdot x+n\] where m is the slope and n is the intercept with the y axis.

OpenStudy (anonymous):

Okay so?

OpenStudy (john_es):

As you have two points, you can form a system of equations. Just plug in the values in the standard form for each point. Do you know how to do it?

OpenStudy (john_es):

For the first porint, (x,y)=(0,5), you have, \[5=m\cdot 0+n\Rightarrow n=5\]

OpenStudy (john_es):

What do you obtain from the second point?

OpenStudy (anonymous):

Ok then?

OpenStudy (john_es):

You should try with the second point (x,y)=(6,0), what do you obtain for m?

OpenStudy (anonymous):

No clue..

OpenStudy (john_es):

Knowing that n=5, you have, \[0=m\cdot 6+5\Rightarrow m=-5/6 \]So the equation you are searching for is, \[y=-\frac{5}{6}x+5\]

OpenStudy (john_es):

These problems can be solved in different forms, but with two points you can solve all of them that way.